OpenStudy (firejay5):

Using Matrices to Solve Systems of Equations and show work!!!! 17. Last Monday at 7:30 a.m., an airline flew 89 passengers on a commuter flight from Boston to New York. Some of the passengers paid $120 for their tickets and the rest paid $230 for their tickets. The total cost of all of the tickets was $14,200. How many passengers bought $120 tickets? How many bought $230 tickets?

OpenStudy (loser66):

Where are you stuck?

OpenStudy (firejay5):

set up of the systems of equations: for example: 4a + 2b = 12

OpenStudy (firejay5):

and matrix part

OpenStudy (loser66):

Nope let a is a number of passenger paid $120 b is a number of passenger paid $230 Then, a +b =89 ok??

OpenStudy (loser66):

"an airline flew 89 passengers....." Make sure that you get the concept. If you get it, let me know to step up

OpenStudy (firejay5):

there are 2 equations

OpenStudy (loser66):

YYYYYYYYYYYes, Sir!! but you have to get the first one first :)

OpenStudy (firejay5):

yes I get it

OpenStudy (loser66):

Now, the second one: then a*120 is the amount of money of the paid-$120 passengers paid b*230 is the amount of money of the paid-$230 passenger paid and 120a + 230b = 14200 get it?

OpenStudy (firejay5):

then since we have the 2 equations we can put them in matrices

OpenStudy (loser66):

yes, I rewrite it: a + b = 89 120a +230b=14200 now, write it as a matrix

OpenStudy (loser66):

\[\left[\begin{matrix}1&1&89\\120&230&14200\end{matrix}\right]\]

OpenStudy (loser66):

get it?

OpenStudy (firejay5):

you don't write 89 and 14200 in the same matrix you write it in a different one. the inverse of matrix A * matrix B

OpenStudy (firejay5):

in a 2x1 matrix = 89 and 14,200

OpenStudy (loser66):

oh!! that way!! oooooooook since I solve it in calculator, I combine them all. But if you have to solve it by that way, it's ok

OpenStudy (firejay5):

we write the answer in an ordered pair

OpenStudy (loser66):

you have: \(\left[\begin{matrix}1&1\\120&230\end{matrix}\right]\)\(\left[\begin{matrix}a\\b\end{matrix}\right]\)=\(\left[\begin{matrix}89\\14200\end{matrix}\right]\)

OpenStudy (loser66):

right?

OpenStudy (firejay5):

yes

OpenStudy (loser66):

so, to find the answer, you need inverse of the first matrix, right?

OpenStudy (firejay5):

right times B and he has the answers in fractions

OpenStudy (loser66):

Ax = B , then x = \(A^{-1}B\), is it what you want?

OpenStudy (firejay5):

yep

OpenStudy (loser66):

ok, do you know how to find inverse of the first matrix?

OpenStudy (loser66):

inverse of this\(\left[\begin{matrix}1&1\\120&230\end{matrix}\right]\)

OpenStudy (loser66):

If you know, try it. When touching matrix, I don't think you can skip any step or rely on others. YOU, have to master it. Try it. I can check. Since you don't know how to combine all, just take step by step, I think you are new in the field. So, Practice, please.

OpenStudy (loser66):

I am waiting for your reply. If you are not interesting on YOUR problem anymore, we stop here. :)

OpenStudy (firejay5):

I think multiply by the matrix 1 0 0 1

OpenStudy (firejay5):

I just click the inverse button on my calculator

OpenStudy (loser66):

nope, if you are allowed to use calculator, then use it to find the inverse of it. Do you know how?

OpenStudy (loser66):

1) enter the whole matrix , not 1 0 0 1

OpenStudy (firejay5):

what's the whole matrix

OpenStudy (loser66):

\(\left[\begin{matrix}1&1\\120&230\end{matrix}\right]\)

OpenStudy (firejay5):

by 1 0 0 1

OpenStudy (loser66):

oh, boy!! ok, again, not 1 0 /0 1

OpenStudy (loser66):

you enter the second matrix: (89 /14200)

OpenStudy (firejay5):

yes I know duh

OpenStudy (loser66):

then hit matrix button to call the first one out, hit the inverse button, DON'T HIT ENTER YET hit times button hit matrix button again to call the second one, then ENTER,

OpenStudy (firejay5):

yep I know how to do it and I got the answer (512/11, 412/11)

OpenStudy (loser66):

What do you get?

OpenStudy (firejay5):

Hello! I got the answer (512/11, 412/11)

OpenStudy (loser66):

NOpe, you must have \(\left[\begin{matrix}57\\32\end{matrix}\right]\) that is the right answer

OpenStudy (loser66):

You make mistake at somewhere. let's think !! a, b are passengers, they must be integers, you can't have 1 and a half person, right? so, if the answer is not integers, you are wrong for sure.